# PDFMathTranslate: Scientific Document Translation Preserving Layouts

### v1

@@ Abstract
Language barriers in scientific documents hinder the diffusion and development of science and technologies. However, prior efforts in translating such documents largely overlooked the information in layouts. To bridge the gap, we introduce PDFMathTranslate, the world’s first open-source software for translating scientific documents while preserving layouts. Leveraging the most recent advances in large language models and precise layout detection, we contribute to the community with key improvements in precision, flexibility, and efficiency. The work has been open-sourced at https://github.com/byaidu/pdfmathtranslate with more than 22k downloads.
@@ 1 Introduction
However, machine translation literature fails to address the unique challenges posed by the layouts in technical translation Schubert (2012). In scientific and technical documents, layouts and non-textual elements are not ignorable — the arrangement of paragraphs, mathematical equations, tables, and figures have rich and important meanings. Those text-based attempts ignore the information in the layout and are thus insufficient to address the barrier in scientific and technical document translation.
@@
To fill the gap, we introduce PDFMathTranslate, the world’s first open-source tool to translate PDF documents with preserved layouts. By leveraging recent advances in layout detection and large language models (see Figure 1), we better address the barrier with at least five key contributions: (1) efficient workflow of layout detection, translation, and re-rendering; (2) support for multiple languages (3) support for multiple translation models and services; (4) diverse user interfaces; and (5) a community-commerce model affording sustainable developments.
@@ 2 Architecture and Design
Overall, PDFMathTranslate is designed to translate documents while preserving their original layouts. First, it accepts a PDF document along with user-specified parameters, such as languages and the preferred translation service (see Appendix A). Next, it detects the layouts to extract the layouts and textual contents from the document. Third, the texts are translated using the selected translation service, such as GPT-4, DeepL, Google, or Ollama (see Appendix A). Finally, the translated text and previously detected layouts are re-rendered as a translated document with preserved layouts.
Technically, we design the architecture for precision, flexibility, and efficiency. Specifically, we introduce three key components: a precise parser, a flexible translation middleware, and an efficient workflow, to be explained in the following three sub-sections.
@@
Preserving layouts prerequires precise parsing. To precisely parse the layouts, we propose a pipeline consisting of layout detection, splitting, processing, and re-rendering. In the beginning, we exploit a recent advance in layout detection,  DocLayout-YOLO-DocStructBench-onnx. The model is another version of a SOTA solution in object detection, Yolov10 Wang et al. , and is fast and accurate in the specific task Zhao et al. .
To increase the compatibility of this model, we support two model formats: the Open Neural Network Exchange (ONNX) onnx standard and a pytorch version. The ONNX version was chosen as default to ensure the compatibility of our parsing pipeline for diverse hardwares.
@@
Translating scientific and technical documents, in some cases, needs an efficient workflow for thousands of pages, challenging the performance. To address the issue, we landed on a streaming design that is centered around an efficient in-memory processing pipeline that minimizes disk I/O while preserving document fidelity. Specifically, the translate_stream function accepts a byte stream of a PDF file and transforms it into a mutable document representation. This approach allows the system to dynamically apply modifications, such as converting to PDF/A format when needed and embedding the appropriate fonts based on the target language. By employing temporary buffers and leveraging functions like download_remote_fonts, the design effectively manages resources and ensures that the original file remains unaltered during processing.
@@ 3 Usage and Deployments
For developers, we support CLI and Docker for advanced usage and substream developments. Developers who prefer command-line tools can conveniently install our Python CLI program using pip install pdf2zh. The command-line tool supports extensive features, with the major ones shown in Figure 3 and documented in detail  [^4] See https://github.com/Byaidu/PDFMathTranslate/blob
@@
. For users aiming to deploy applications on servers, we offer a Docker image designed specifically to simplify online production deployment. Those images are built and distributed on different platforms for developers with potential network issues.
@@
Our comparison indicates that PDFMathTranslate surpasses alternative solutions. First, regarding accessibility, we significantly reduce user costs by providing an open-source, self-deployable, and free solution with flexible API integration. This flexibility enables broad applicability of our tool and ensures accurate translation outputs from any preferred service or model.Second, in terms of readability, our approach preserves basic layouts, accurately recognizes complex mathematical formulas, and supports bilingual outputs. Finally, although our solution is not the fastest—due to the computational demands of on-device inference—its efficiency remains comparable to other services (see 1).These advantages suggest that our tool becomes a better alternative to existing tools in the translation of scientific documents, thus potentially better reducing the language barriers.
@@ 4 Use Cases
We illustrate two typical use cases of PDFMathTranslate to illustrate how the tool translates text while preserving information embedded in layouts. The first use case is a technical document with both text and complex formulas, and the second use case is a research article with complex layouts and figures (see Figure 4).
Both cases show that the tool is capable of precisely translating the information within text while preserving the crucial information embedded in layouts.
@@
5 Comparison with Existing Systems
To illustrate the advantages of our tool over other commercial and open-source alternatives in several important aspects, including accessibility, readability, and efficiency.
@@ 6 Sustainability
Nevertheless, maintenance is challenging in open-source projects Stol and Ali Babar . To ensure sustainable development, we’ve established a community-commerce model, through which the incentives of developers are provided by two sources: open-source recognition and the benefits of commercial products.
@@
Second, through collaboration with a commercial partner, we provide sponsored rewards such as membership exchanges for active contributors. These 11 subscriptions sent out to developers have effectively encouraged consistent contributions, resulting in the resolution of over 485 user-reported issues (by April 2025). Such cooperation has enabled the integration of advanced academic and community-driven technologies, including ONNX model support Wybxc (2025) and Qwen ws051682 (2025), thus advancing cross-linguistic scientific communication.
This sustainable collaboration model propelled the project to the top of GitHub’s global trends for over a week, garnering more than 25k stars, 22k downloads, and over 49k Docker pulls (by June 2025). Additionally, our project’s success has inspired a range of new commercial products with similar functionalities, significantly impacting both the scientific community and the broader public.
@@ 7 Limitations
While we have made several key improvements, our tool still has specific limitations. First, translation quality remains highly dependent on the underlying translation models and prompts. Second, the accuracy of the layout detection model in identifying various document layouts similarly depends on the quality of the layout detection models employed. Third, the current version of the tool can not yet handle scanned PDFs lacking optimal optical character recognition (OCR); however, this last feature is planned for implementation.
@@ Appendix A Appendix
A.1 Appendix 1. Language and services supported in PDFMathTranslate
@@
Service or Google (default), Bing, DeepL, DeepLX, Ollama, Ali Qwen Translation, 22 a
models Ollama Xinference, Gemma, OpenAI, OpenAI-likea, DeepSeek,
@@
Input and Amharic, Arabic, Basque, Bengali, English (UK), 56 b
@@
  • a
    Note. a The number of supported languages far exceeds what is shown in the table. Our protocol, designed as a general framework for OpenAI-like services, allows users to integrate any translation system, including locally hosted models or private online services, ensuring extensive language support.
  • b
    b Based on the default service (Google Translate) and interface (CLI). However, the actual number may be significantly higher due to the extensibility of services and models.

### v2

@@ Abstract
Language barriers in scientific documents hinder the diffusion and development of science and technologies. However, prior efforts in translating such documents largely overlooked the information in layouts. To bridge the gap, we introduce PDFMathTranslate, the world’s first open-source software for translating scientific documents while preserving layouts. Leveraging the most recent advances in large language models and precise layout detection, we contribute to the community with key improvements in precision, flexibility, and efficiency. The work is open-sourced at https://github.com/byaidu/pdfmathtranslate with more than 222k downloads.
@@ 1 Introduction
However, machine translation literature fails to address the unique challenges posed by the layouts in technical translation Schubert (2012). In scientific and technical documents, layouts and non-textual elements are not ignorable — the arrangement of paragraphs, mathematical equations, tables, and figures has rich and important meanings. Those text-based attempts ignore the information in the layout and are thus insufficient to address the barrier in scientific and technical document translation.
@@
To fill the gap, we introduce PDFMathTranslate, the world’s first open-source tool to translate PDF documents with preserved layouts. By leveraging recent advances in layout detection and large language models (see Figure 1), we better address the barrier with at least five key contributions: (1) efficient workflow of layout detection, translation, and re-rendering; (2) support for multiple languages; (3) support for multiple translation models and services; (4) diverse user interfaces; and (5) a community-commerce model affording sustainable developments.
@@ 2 Architecture and Design
PDFMathTranslate is designed to translate documents while preserving their original layouts. First, it accepts a PDF document along with user-specified parameters, such as languages and the preferred translation service (see Appendix A). Next, it detects the layouts to extract the layouts and textual contents from the document. Third, the texts are translated using the selected translation service, such as GPT-4, DeepL, Google, or Ollama (see Appendix A). Finally, the translated text and previously detected layouts are re-rendered as a translated document with preserved layouts.
Technically, we design the architecture for precision, flexibility, and efficiency. We introduce three key components: a precise parser, a flexible translation middleware, and an efficient workflow, which we will explain in the following three subsections.
@@
Preserving layouts requires precise layout extraction. To precisely parse the position of document elements, we propose a pipeline consisting of layout detection, splitting, processing, and re-rendering. In the beginning, we exploit a recent advance in layout detection,  DocLayout-YOLO-DocStructBench-onnx. The model is another version of a SOTA solution in object detection, Yolov10 Wang et al. , and is fast and accurate in the specific task Zhao et al. .
To increase the compatibility of this model, we support two model formats: the Open Neural Network Exchange (ONNX) onnx standard and a pytorch version. The ONNX version was chosen as default to ensure the compatibility of our parsing pipeline for diverse hardware.
@@
Translating scientific and technical documents often necessitates an efficient workflow, especially when dealing with thousands of pages, which can challenge overall performance. To address the issue, we landed on a streaming design that is centered around an efficient in-memory processing pipeline that minimizes disk I/O while preserving document fidelity. Specifically, the translate_stream function accepts a byte stream of a PDF file and transforms it into a mutable document representation. This approach allows the system to dynamically apply modifications, such as converting to PDF/A format when needed and embedding the appropriate fonts based on the target language. By employing temporary buffers and leveraging functions like download_remote_fonts, the design effectively manages resources and ensures that the original file remains unaltered during processing.
@@ 3 Usage and Deployments
For developers, we support CLI and Docker for advanced usage and substream developments. Developers who prefer command-line tools can conveniently install our Python CLI program using pip install pdf2zh. The command-line tool supports extensive features, with the major options shown in Figure 3 and more documented in details  [^4] See https://github.com/Byaidu/PDFMathTranslate/blob
@@
. For users who want to deploy the software on servers, we offer a Docker image for simplified deployment. Additionally, Docker images are built and distributed on different platforms for developers with potential network issues.
@@
4 Comparison with Existing Systems
Compared with existing alternatives, PDFMathTranslate offers a superior solution from three perspectives: accessibility, readability, and efficiency (see Table 1).
In terms of accessibility, the tool is open-sourced, self-deployable, supports API calls, and is entirely free of charge. These advantages ensure wider accessibility compared to any other existing solutions. Regarding readability, the tool preserves layouts, supports complex formulas, and is capable of exporting bilingual documents. Notably, a current limitation of our tool is its lack of optimal optical character recognition (OCR) support, which restricts its wider application for scanned documents. However, OCR functionality is planned for implementation in upcoming versions. Finally, concerning efficiency, while this application is slower than text-based translation services (like Google Translate) due to the integration of precise layout detection models, our solution is significantly faster than alternative products that also preserve layouts (such as IMT and Doc2X).
@@ 5 Use Cases
We illustrate two typical use cases of PDFMathTranslate to illustrate how the tool translates text while preserving information embedded in layouts. The first use case (Panel A, Figure 4) is an excerpt from a textbook with both text and complex formulas, and the second case (Panel B, Figure 4) is a scientific research article with complex layouts and figures (see Figure 4). Both cases consistently show that our work is capable of precisely translating the information within text while preserving the crucial information embedded in layouts.
@@ 6 Sustainability
Nevertheless, maintenance is challenging to open-source projects Stol and Ali Babar . To ensure sustainable development, we’ve established a community-commerce model, through which the incentives of developers are provided by two sources: open-source recognition and the benefits of commercial products.
@@
Second, we provide sponsored rewards such as membership exchanges for active contributors through collaboration with a commercial partner. These 11 subscriptions sent out to developers have effectively encouraged consistent contributions, resulting in the resolution of over 485 user-reported issues (by April 2025). Such cooperation has enabled the integration of advanced academic and community-driven technologies, including ONNX model support Wybxc (2025) and Qwen ws051682 (2025), thus advancing cross-linguistic scientific communication.
This sustainable collaboration model propelled the project to the top of GitHub’s global trends for over a week, garnering more than 25k stars, 222k downloads, and over 49k Docker pulls (by June 2025). Additionally, our project’s success has inspired a range of new commercial products with similar functionalities, significantly impacting both the scientific community and the broader public.
@@ 7 Limitations
While we have made several key improvements, our tool still has specific limitations. First, translation quality remains highly dependent on the underlying translation models and prompts. Second, the accuracy of the layout detection model in identifying various document layouts similarly depends on the quality of the layout detection models employed. Third, the current version of the tool can not handle scanned PDFs lacking optimal optical character recognition (OCR); however, this last feature is planned for implementation.
@@ Appendix A Appendix
A.1 Officially supported language and services
@@
Service or Google (default), Bing, DeepL, DeepLX, Ollama, Ali Qwen Translation, 22 1
models Ollama Xinference, Gemma, OpenAI, OpenAI-like 1, DeepSeek,
@@
Input and Amharic, Arabic, Basque, Bengali, English (UK), 56 2
@@
  • 1
    1 The number of supported languages far exceeds what is shown in the table. Our protocol, designed as a general framework for OpenAI-like services, allows users to integrate any translation system, including locally hosted models or private online services, ensuring extensive language support.
  • 2
    2 Based on the default service (Google Translate) and interface (CLI). However, the actual number may be significantly higher due to the extensibility of services and models.